OpenStudy (anonymous):

what is the slope of (4,5) and (6,9) ?

OpenStudy (anonymous):

4/2 = 2 the slope is 2

OpenStudy (anonymous):

the slope is 2 hope i helpedd if needed explanation just ask!

OpenStudy (anonymous):

take the difference of each coordinate (x2-x1)=(6-4)=2 (y2-y1)=(9-5)=4 then do rise over run 4/2=2